# A unified concept of the degree of ill-posedness for compact and non-compact linear operators in Hilbert spaces under the auspices of the spectral theorem

### v1

@@ 5 The case of unbounded operators
As in Section 3 we discuss the application of the spectral theorem, but with one substantial change: the injective linear operator $T:D(T)\subset\mathcal{X}\to\mathcal{Y}$ with domain $D(T)$ dense in $\mathcal{X}$ and hence its self-adjoint non-negative companion $T^{*}T:D(T)\subset\mathcal{X}\to\mathcal{X}$ with $0\in{\rm spec}(T^{*}T)$ are unbounded and consequently not compact. Then the spectral theorem in multiplication operator version is also available, and we refer for details to [35, Chapt. VII.3]. In particular, now in the factorization (5) there occurs an unbounded multiplication operator $M_{\lambda}$ (cf. (6)) with a $\mu$-measurable multiplier function $\lambda:\Omega\to\mathbb{R}$ such that $\min{\rm essran}(\lambda)=0$ and $\sup{\rm essran}(\lambda)=+\infty$, since we have again the coincidence ${\rm spec}(T^{*}T)={\rm essran}(\lambda)$. Let Assumption 1 be valid, which implies that $\mu(\Omega)=\infty$, the distribution function $\Phi_{\lambda}(\epsilon)$ (cf. (7)) is finite for all $\epsilon>0$ and its inverse $\lambda^{*}$, the decreasing rearrangement of $\lambda$, is well-defined and an index function at the infinity. It is not difficult to understand that poles of $\lambda$ do not really influence the decay rate of $\lambda^{*}(t)$ as $t\to\infty$. Thus, Definition 2 seems to be applicable also here for verifying the degree of ill-posedness. Regularization approaches for ill-posed operator equations with unbounded forward operators are, for example, discussed in [18] and [34] by exploiting the role of corresponding multiplier functions.
@@
One easily finds that here ${\rm spec}(T^{*}T)={\rm essran}(\lambda)=[0,\infty)$, which means that $T^{*}T$ is a non-compact, unbounded linear operator, because $\lambda(\omega)$ has a pole at $\omega=0$. Moreover, ill-posedness takes place with $\lim_{\|\omega\|_{2}\to\infty}\lambda(\omega)=0$. Now we have $\lambda(\omega)\asymp\|\omega\|_{2}^{-4}$ as $\|\omega\|_{2}\to\infty$, which should imply for the decreasing rearrangement $\lambda^{*}$ of $\lambda$ that $\lambda^{*}(t)\asymp t^{-\frac{4}{d}}$ as $t\to\infty$, which indicates moderate ill-posedness with degree $s=\frac{2}{d}$. By applying for sufficiently small $\epsilon>0$ the distribution function $\Phi_{\lambda}(\epsilon)\approx\mu_{\mathrm{Leb}}(\{\omega\in\mathbb{R}^{d}:\|\omega\|_{2}\leq\frac{1}{\kappa\epsilon^{1/4})}\})$, the denominator $d$ in the asymptotics of $\lambda^{*}$ arises from the fact that the volume of a ball in $\mathbb{R}^{d}$ with radius $\frac{1}{\kappa\epsilon^{1/4}}$ is proportional to $\epsilon^{-d/4}$.
We refer to [34] for more details and the discussion of the full parabolic source identification problem in the multidimensional case. Let us note here that such degree of ill-posedness $s=\frac{2}{d}$ with dimension $d$ in the denominator also occurs for various compact multivariate forward operators, e.g. for the inverse of the negative Laplacian in $L^{2}(\Sigma)$ with homogeneous boundary conditions on a bounded domain $\Sigma\subset\mathbb{R}^{d}$ and with such $\Sigma$ for the embedding operator from $H^{1}(\Sigma)$ into $L^{2}(\Sigma)$.
@@ 6 Conclusion and outlook
In this paper, we have presented a unified framework for measuring the degree of ill-posedness for linear operators by means of its spectral decomposition. We have shown that the derived concept coincides with the well-known definition in the compact case and allows us to interpret several non-compact operators as moderately or severely ill-posed as well.
However, currently our concept is not able to handle all bounded linear operators. The special situations (A), (B) and (C) in Figure 1 have to be excluded for different reasons. It will be an interesting topic for future research to investigate whether a situation of finite measure $\mu(\Omega)$ can be handled similarly or not. Furthermore it is unclear if other concepts to measure ill-posedness by means of essential zeros of $\lambda$ in the situation (B) can be related to our concept in some way. Finally, it would be an interesting question to find real-world examples for the situation (C). This mostly seems to be a measure-theoretic problem.
@@ Acknowledgment
In this appendix, we return to the case of bounded linear operators $T:\mathcal{X}\to\mathcal{Y}$, and we will briefly summarize references and specific approaches for the alternative case of finite measures $\mu(\Omega)<\infty$. This case has been comprehensively studied in the literature for non-compact multiplication operators $T$ mapping in $L^{2}(\Omega)$ for some bounded subinterval $\Omega$ of $\mathbb{R}$. The focus of most papers is on the unit interval $\Omega=[0,1]$ and the Lebesgue measure $\mu=\mu_{\mathrm{Leb}}$ over $\mathbb{R}$. For that measure, non-negative multiplier functions $\lambda\in L^{\infty}(0,1)$ are under consideration that possess essential zeros inside $[0,1]$. This leads to ill-posed situations of the corresponding operator equations (1), where the structure of the zeros seem to play a prominent role for the strength of the ill-posedness. We refer in this context to [7, 13, 15, 20, 21] and recently [26]. In the latter reference, in particular, the total influence of different essential zeros of the multiplier function $\lambda$ on the increasing rearrangement of $\lambda$ is discussed. This increasing rearrangement and its growth rate at a neighbourhood of zero are tools for characterizing the degree of ill-posedness for this class of problems with finite measure spaces.
Unfortunately, the cross connections of this increasing rearrangement concept for $\Omega=[0,1]$ and the associated decreasing rearrangement concept for $\mu(\Omega)=\infty$ mentioned in Remark 4 above are not completely clear. It is appropriate for the finite measure case to use instead of (7) the non-decreasing distribution function
@@
which is the (generalized) inverse to the function $d_{\lambda}$. Both $d_{\lambda}$ and $\lambda^{*}$ are index function at zero, which means that they are non-decreasing functions with positive values for positive arguments and zero limits as the arguments tend to zero. Then the most authors characterize the ill-posedness by the growth rate of the function $\lambda^{*}(t)$ in a right neighborhood of $t=0$. As shown in [26, Proposition 2] the strongest essential zero of the multiplier $\lambda$ over the interval $[0,1]$ is responsible for the degree of ill-posedness.

### v2

@@ 5 The case of unbounded operators
As in Section 3 we discuss the application of the spectral theorem, but with one substantial change: the injective linear operator $T:D(T)\subset\mathcal{X}\to\mathcal{Y}$ with domain $D(T)$ dense in $\mathcal{X}$ and hence its self-adjoint non-negative companion $T^{*}T:D(T)\subset\mathcal{X}\to\mathcal{X}$ with $0\in{\rm spec}(T^{*}T)$ are unbounded and consequently not compact. Then the spectral theorem in multiplication operator version is also available, and we refer for details to [37, Chapt. VII.3]. In particular, now in the factorization (5) there occurs an unbounded multiplication operator $M_{\lambda}$ (cf. (6)) with a $\mu$-measurable multiplier function $\lambda:\Omega\to\mathbb{R}$ such that $\min{\rm essran}(\lambda)=0$ and $\sup{\rm essran}(\lambda)=+\infty$, since we have again the coincidence ${\rm spec}(T^{*}T)={\rm essran}(\lambda)$. Let Assumption 1 be valid, which implies that $\mu(\Omega)=\infty$, the distribution function $\Phi_{\lambda,\mu_{Leb}}(\epsilon)$ (cf. (7)) is finite for all $\epsilon>0$ and its inverse $\lambda^{*}$, the decreasing rearrangement of $\lambda$, is well-defined and an index function at the infinity. It is not difficult to understand that poles of $\lambda$ do not really influence the decay rate of $\lambda^{*}(t)$ as $t\to\infty$. Thus, Definition 2 seems to be applicable also here for verifying the degree of ill-posedness. Regularization approaches for ill-posed operator equations with unbounded forward operators are, for example, discussed in [19] and [36] by exploiting the role of corresponding multiplier functions.
@@
One easily finds that here ${\rm spec}(T^{*}T)={\rm essran}(\lambda)=[0,\infty)$, which means that $T^{*}T$ is a non-compact, unbounded linear operator, because $\lambda(\omega)$ has a pole at $\omega=0$. Moreover, ill-posedness takes place with $\lim_{\|\omega\|_{2}\to\infty}\lambda(\omega)=0$. Now we have $\lambda(\omega)\asymp\|\omega\|_{2}^{-4}$ as $\|\omega\|_{2}\to\infty$, which should imply for the decreasing rearrangement $\lambda^{*}$ of $\lambda$ that $\lambda^{*}(t)\asymp t^{-\frac{4}{d}}$ as $t\to\infty$, which indicates moderate ill-posedness with degree $s=\frac{2}{d}$. By applying for sufficiently small $\epsilon>0$ the distribution function $\Phi_{\lambda,\mu_{Leb}}(\epsilon)\approx\mu_{\mathrm{Leb}}(\{\omega\in\mathbb{R}^{d}:\|\omega\|_{2}\leq\frac{1}{\kappa\epsilon^{1/4})}\})$, the denominator $d$ in the asymptotics of $\lambda^{*}$ arises from the fact that the volume of a ball in $\mathbb{R}^{d}$ with radius $\frac{1}{\kappa\epsilon^{1/4}}$ is proportional to $\epsilon^{-d/4}$.
We refer to [36] for more details and the discussion of the full parabolic source identification problem in the multidimensional case. Let us note here that such degree of ill-posedness $s=\frac{2}{d}$ with dimension $d$ in the denominator also occurs for various compact multivariate forward operators, e.g. for the inverse of the negative Laplacian in $L^{2}(\Sigma)$ with homogeneous boundary conditions on a bounded domain $\Sigma\subset\mathbb{R}^{d}$ and with such $\Sigma$ for the embedding operator from $H^{1}(\Sigma)$ into $L^{2}(\Sigma)$ (see also Example 3).
@@ 6 Conclusion and outlook
In this paper, we have presented a unified framework for measuring the degree of ill-posedness for linear operator equations by means of its spectral decomposition. We have shown that the derived concept coincides with the well-known definition in the compact case and allows us to interpret several non-compact operator equations as moderately or severely ill-posed as well when the Lebesgue measure is chosen as benchmark case. Unfortunately, cross comparisons between compact and non-compact operators cannot be handled by our framework, because compact operators and the Lebesgue measure are not compatible with respect to multiplication operators, and we refer to Proposition 2 in this context.
Currently our concept is also limited concerning the special situations (A), (B) and (C) in Figure 1, which have to be excluded for different reasons. It will be an interesting topic for future research to investigate whether a situation of finite measure $\mu(\Omega)$ can be handled similarly or not. Furthermore it is unclear if other concepts to measure ill-posedness by means of essential zeros of $\lambda$ in the situation (B) can be related to our concept in some way. Finally, it would be an interesting question to find real-world examples for the situation (C). This mostly seems to be a measure-theoretic problem.
@@ Acknowledgment
In this appendix, we return to the case of bounded linear operators $T:\mathcal{X}\to\mathcal{Y}$, and we will briefly summarize references and specific approaches for the alternative case of finite measures $\mu(\Omega)<\infty$. This case has been comprehensively studied in the literature for non-compact multiplication operators $T$ mapping in $L^{2}(\Omega)$ for some bounded subinterval $\Omega$ of $\mathbb{R}$. The focus of most papers is on the unit interval $\Omega=[0,1]$ and the Lebesgue measure $\mu=\mu_{\mathrm{Leb}}$ over $\mathbb{R}$. For that measure, non-negative multiplier functions $\lambda\in L^{\infty}(0,1)$ are under consideration that possess essential zeros inside $[0,1]$. This leads to ill-posed situations of the corresponding operator equations (1), where the structure of the zeros seem to play a prominent role for the strength of the ill-posedness. We refer in this context to [8, 14, 16, 21, 22] and recently [28]. In the latter reference, in particular, the total influence of different essential zeros of the multiplier function $\lambda$ on the increasing rearrangement of $\lambda$ is discussed. This increasing rearrangement (see for details [7]) and its growth rate at a neighbourhood of zero are tools for characterizing the degree of ill-posedness for this class of non-compact operators modelled with finite measure spaces.
Unfortunately, the cross connections of this increasing rearrangement concept for $\Omega=[0,1]$ and the associated decreasing rearrangement concept for $\mu(\Omega)=\infty$, see formula (10) above, are not completely clear. It is appropriate for the finite measure case to use instead of (7) the non-decreasing distribution function
@@
which is the (generalized) inverse to the function $d_{\lambda}$. Both $d_{\lambda}$ and $\lambda^{*}$ are index functions at zero, which means that they are non-decreasing functions with positive values for positive arguments and zero limits as the arguments tend to zero. Then the most authors characterize the ill-posedness by the growth rate of the function $\lambda^{*}(t)$ in a right neighborhood of $t=0$. As shown in [28, Proposition 2] the strongest essential zero of the multiplier $\lambda$ over the interval $[0,1]$ is responsible for the degree of ill-posedness.